# BiDoRA: Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation

### v1

@@ 4.1 Overview of BiDoRA
Our method, BiDoRA, optimizes the trainable parameters in DoRA layers by solving a bi-level optimization problem. Let =$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$ denote the set of magnitude components for all $n$ DoRA modules, and $\mathcal{V}=\{\Delta V_{1},\Delta V_{2},\ldots,\Delta V_{n}\}$ denote the set of corresponding incremental direction components. Specifically, we first learn the incremental direction components $\mathcal{V}^{*}(\mathcal{M})$ on the training split of the downstream dataset $\mathcal{D}_{tr}$ at the lower level. The magnitude component $\mathcal{M}$ is tentatively fixed at this level, thus the resulting optimal incremental direction component $\mathcal{V}^{*}(\mathcal{M})$ is a function of $\mathcal{M}$. At the upper level, we determine the optimal magnitude component $\mathcal{M}^{*}$ by optimizing the loss on a validation split $\mathcal{D}_{val}$. In practice, $\mathcal{D}_{tr}$ and $\mathcal{D}_{val}$ are typically created by splitting the original training set without using additional data. This bi-level optimization problem is solved using an efficient gradient-based algorithm, where parameters in two levels are optimized iteratively until convergence.
@@ 4.2 Orthogonal Regularization
The orthogonality of neural network weights has been identified as a beneficial property (Bansal et al., 2018) and can effectively mitigate the overfitting issue (Balestriero & richard baraniuk, 2018). Therefore, we define a Gram regularization loss (Xie et al., 2017) for the direction component:
$\mathcal{R}(\mathcal{V})=\sum_{k=1}^{n}\left\lVert(V_{k}+\Delta V_{k})^{\top}(V_{k}+\Delta V_{k})-I\right\rVert_{F}^{2}$ (3)
where $I$ is the identity matrix and $\|\cdot\|_{F}$ denotes the Frobenius norm. Intuitively, $\mathcal{R}(\mathcal{V})$ encourages each column of the direction matrix, representing a specific direction, to be orthogonal to one another. Since each column has already been normalized (equivalent to projected to the unit sphere), this also prompts each column to be far away from the other, thereby reducing the redundancy of parameters.
@@ 4.3 A Bi-level Optimization Framework
Lower Level
At the lower level, we train the low-rank incremental direction component $\mathcal{V}$ by minimizing a loss $\mathcal{L}_{tr}$ defined on the training set $\mathcal{D}_{tr}$. The overall training objective at this level is $\mathcal{L}_{tr}(\mathcal{V},\mathcal{M})=C(\mathcal{V},\mathcal{M};\mathcal{D}_{tr})+\gamma\mathcal{R}(\mathcal{V})$. Here, $C$ represents the fine-tuning loss, given the low-rank incremental direction component $\mathcal{V}$, the magnitude component $\mathcal{M}$, and the training split $\mathcal{D}_{tr}$ of the downstream dataset. $\mathcal{R}(\mathcal{V})$ is the orthogonal regularizer defined in Eq. (3), with $\gamma$ as a trade-off hyperparameter. In this level, we only update $\mathcal{V}$ while keeping $\mathcal{M}$ fixed, resulting in the following optimization problem:
@@
Upper Level
At the upper level, we validate the previously fixed magnitudes $\mathcal{M}$ on the validation set $\mathcal{D}_{val}$, using the optimal incremental direction component $\mathcal{V}^{*}(\mathcal{M})$ that was learned at the lower level. This results in a validation loss $\mathcal{L}_{val}(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M})=C(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M};\mathcal{D}_{val})$. We determine the optimal magnitude component $\mathcal{M}$ by minimizing this validation loss:
@@
A Bi-level Optimization Framework
Integrating the two levels of optimization problems, we have the following bi-level optimization framework:
@@ A Bi-level Optimization Framework
Two reasons exist behind the choice of setting the magnitude component as the upper level instead of the converse one: 1) In literature, the upper level usually has much fewer parameters than the lower level. In our case, the design of setting the magnitude of complexity $\mathcal{O}(k)$ as the upper level and the direction of complexity $\mathcal{O}(dr+kr)$ as the lower level is consistent with the common practice. 2) BiDoRA resembles the DARTS method (Liu et al., 2018) in neural architecture search where the subnets are selected by a selection variable. Specifically, the magnitude vector resembles a selection variable on the direction matrix by softly selecting each direction (subnets) via scaling.
Optimization Algorithm
@@
2Initialize trainable magnitude components =$\mathcal{M}=\{m_{k}\}_{k=1}^{n}$ and low-rank incremental direction components $\mathcal{V}=\{\Delta V_{k}\}_{k=1}^{n}=\{\{A_{k}\}_{k=1}^{n},\{B_{k}\}_{k=1}^{n}\}$
@@
We use a gradient-based optimization algorithm (Choe et al., 2023) to solve the bi-level optimization problem presented in Eq. (4.3). A significant challenge in this process is that precisely computing the gradient of the upper-level loss $\mathcal{L}_{val}$ with respect to the magnitude component $\mathcal{M}$ can be computationally prohibitive due to the lack of an analytical solution for $\mathcal{V}^{*}(\mathcal{M})$ at the lower-level optimization problem. To address this issue, we use the following one-step-unrolled approximation of $\mathcal{V}^{*}(\mathcal{M})$ inspired by previous work (Liu et al., 2018):
$\nabla_{\mathcal{M}}{\mathcal{L}_{val}(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M})}\approx\nabla_{\mathcal{M}}{\mathcal{L}_{val}(\mathcal{V}-\xi\nabla_{\mathcal{V}}{\mathcal{L}_{tr}(\mathcal{V},\mathcal{M})},\mathcal{M})}$
@@
where $\epsilon$ is a small scalar and $\mathcal{V}^{\pm}=\mathcal{V}\pm\epsilon\nabla_{\bar{\mathcal{V}}}{\mathcal{L}_{val}(\bar{\mathcal{V}},\mathcal{M})}$. Since directly computing the matrix-vector multiplication term in Eq. (7) is computationally expensive, we use finite difference to approximate this product as in Eq. (8), following Liu et al. (2018). As detailed in Algorithm 1, the incremental direction component $\mathcal{V}$ and the magnitude component $\mathcal{M}$ are updated using gradient descent iteratively until convergence. After acquiring the optimal magnitudes $\mathcal{M}^{*}$ through the process above, the incremental direction component $\mathcal{V}$ is retrained on the union of training and validation splits to achieve the best performance on downstream tasks, resulting in the final learned $\mathcal{V}^{*}$.
@@ 5.1 Experimental Setup
We compare BiDoRA with several PEFT methods, including Adapter tuning (Houlsby et al., 2019), LoRA (Hu et al., 2021), and DoRA (Liu et al., 2024). BiDoRA does not use any additional data compared to other baselines, as we create the validation set for upper-level optimization by splitting the original training set with an $8$:$2$ ratio for all tasks. All methods in the experiment, including ablation studies, are trained until convergence for a fair comparison. Detailed descriptions of these baseline methods are provided in Appendix C.
Table 1: RoBERTabase/large (Rb/l) with different fine-tuning methods on the GLUE benchmark. A higher value is better for all datasets. The best results are shown in bold.
Method Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg
Rb(FT) $125.0$M $90.3$ $94.8$ $89.3$ $61.6$ $86.7$ $92.8$ $76.9$ $91.2$ $85.5$
Rb(Adapter) $0.9$ M $86.5$ $94.0$ $88.4$ $58.8$ $92.5$ $89.1$ $71.2$ $89.9$ $83.8$
Rb(LoRA) $0.15$ M $86.8$ $94.3$ $88.0$ $60.3$ 93.0 $89.6$ $72.9$ $90.1$ $84.4$
Rb(DoRA) $0.17$ M $86.8$ $94.2$ $89.2$ $60.5$ $92.9$ $89.6$ $73.2$ 90.2 $84.6$
Rb(BiDoRA) $0.17$ M 87.1 94.4 89.4 61.3 $92.7$ $90.6$ 76.0 $90.1$ 85.2
Rl(FT) $355.0$M $90.2$ $96.4$ $90.9$ $68.0$ $94.7$ $92.2$ $86.6$ $92.4$ $88.9$
Rl(Adapter) $0.8$M $90.3$ $96.3$ $87.7$ $66.3$ 94.7 $91.5$ $72.9$ $91.5$ $86.4$
Rl(LoRA) $0.39$ M 90.6 $96.3$ $90.0$ $66.9$ $94.5$ $91.2$ $86.3$ $91.7$ $88.4$
Rl(DoRA) $0.39$ M 90.6 96.4 $89.8$ $65.8$ 94.7 $91.2$ $86.6$ 92.0 $88.4$
Rl(BiDoRA) $0.39$ M 90.6 $96.1$ 90.1 67.0 $94.6$ 91.7 86.9 92.0 88.6
Table 2: RoBERTabase/large (Rb/l) with different fine-tuning methods on the Reuters21578 benchmark. A higher value is better for all datasets. The best results are shown in bold.
Method Param ModApte ModHayes ModLewis
Rb(FT) $125.0$M $85.4$ $77.6$ $77.1$
Rb(Adapter) $0.9$ M 85.3 $77.5$ $76.8$
Rb(LoRA) $0.15$ M $84.7$ $74.3$ $74.7$
Rb(DoRA) $0.17$ M $84.8$ $78.2$ $76.6$
Rb(BiDoRA) $0.17$ M 85.3 79.9 77.6
Rl(FT) $355.0$M $84.8$ $77.5$ $76.6$
Rl(Adapter) $0.44$ M $84.8$ $77.9$ $76.7$
Rl(LoRA) $0.39$ M $84.7$ $77.7$ $76.7$
Rl(DoRA) $0.39$ M $84.8$ $77.4$ $76.7$
Rl(BiDoRA) $0.39$ M 84.9 78.9 77.3
Our experiments covere a wide range of tasks, including natural language understanding (NLU), natural language generation (NLG), and token classification. For NLU tasks, we fine-tune the RoBERTa-base and RoBERTa-large models on the GLUE benchmark (Wang et al., 2019) and the Reuters21578 dataset (Padmanabhan et al., 2016) using all baseline PEFT methods and BiDoRA. Detailed descriptions of these datasets and pre-trained models are provided in Appendix A. Following existing practices, the development set is used in GLUE as the test data since the actual test set is not publicly available. We report the overall (matched and mismatched) accuracy for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, and accuracy for the other tasks. On the Reuters21578 dataset, the F1 score is used as the evaluation metric across all three splits. For NLG tasks, we fine-tune GPT-2 medium on the E2E (Novikova et al., 2017) dataset. We use BLEU (Papineni et al., 2002), NIST (Lin & Och, 2004), METEOR (Banerjee & Lavie, 2005), ROUGE-L (Lin, 2004), and CIDEr (Vedantam et al., 2015) as evaluation metrics. For token classification, we fine-tune the RoBERTa-base and RoBERTa-large models on the BioNLP (Collier et al., 2004) dataset and the CoNLL2003 (Sang & De Meulder, 2003) dataset. Accuracy, precision, recall, and F1 score are used as evaluation metrics.
For all experiments, our implementation is based on the Huggingface Transformers library (Wolf et al., 2019) and the Betty library (Choe et al., 2023). We use a single NVIDIA A100 GPU for all experiments. More detailed experimental settings are provided in Appendix B.
@@ 5.2 Experiments on Natural Language Understanding Tasks
In this section, we evaluate the performance of BiDoRA on NLU tasks, with a particular focus on text classification. Table 1 presents the results of fine-tuning the RoBERTa-base and RoBERTa-large models on the GLUE benchmark with baseline PEFT methods and BiDoRA. The results show that BiDoRA achieves superior or comparable performance compared to baseline methods across all datasets with the same number of trainable parameters. Table 2 presents the results of fine-tuning RoBERTa models on the Reuters21578 datasets, where BiDoRA outperforms all baseline methods by an even larger margin. Notably, BiDoRA achieves performance comparable to or even better than full fine-tuning. The superior performance of BiDoRA on both benchmarks verifies the effectiveness of its bi-level optimization mechanism. By training the magnitude and incremental direction components on two distinct sub-datasets, BiDoRA enhances the flexibility of the learning process and improves learning capacity compared to DoRA, resulting in a performance boost.
Table 3: Performance of BiDoRA and baseline methods for fine-tuning GPT2-medium on the E2E dataset. A higher value is better for all metrics. The best results are shown in bold.
Method Param BLEU NIST MET ROUGE-L CIDEr
Full FT $354.9$M $68.0$ $8.61$ $46.1$ $69.0$ $2.38$
Adapter $11.1$M $67.0$ $8.50$ $45.2$ $66.9$ $2.31$
LoRA $0.39$M $63.7$ $8.19$ $44.1$ $68.0$ $2.17$
DoRA $0.39$M $67.0$ $8.48$ $45.4$ $70.1$ $2.33$
BiDoRA $0.39$M 69.0 8.72 46.2 70.9 2.44
5.3 Experiments on Natural Language Generation Tasks
In this section, we evaluate BiDoRA’s performance on the NLG task. Table 3 presents the results of fine-tuning a GPT-2 model on the E2E dataset with baseline PEFT methods and BiDoRA. The results show that BiDoRA achieves the best performance across all five evaluation metrics, demonstrating the superiority of BiDoRA in fine-tuning pre-trained models for NLG tasks.
5.4 Experiments on Token Classification
Table 4: RoBERTabase/large (Rb/l) with different fine-tuning methods on BioNLP data and CoNLL2003 dataset. A higher value is better for all metrics. The best results are shown in bold.
BioNLP CoNLL2003
Method Param Accuracy Precision Recall F1 Accuracy Precision Recall F1
Rb(FT) $125.0$M $93.9$ $69.0$ $78.9$ $73.6$ $99.3$ $95.7$ $96.3$ $96.0$
Rb(Adapter) $0.9$ M $93.9$ $69.1$ $78.8$ $73.7$ 99.3 $95.7$ $96.4$ $96.0$
Rb(LoRA) $0.15$ M $93.9$ $69.0$ $78.8$ $73.6$ 99.3 $95.4$ $96.3$ $95.8$
Rb(DoRA) $0.17$ M 94.0 $69.2$ 79.1 $73.8$ 99.3 $95.3$ $96.2$ $95.8$
Rb(BiDoRA) $0.17$ M $93.9$ 71.2 $78.6$ 74.7 99.3 95.9 96.5 96.2
Rl(FT) $355.0$M $94.0$ $69.4$ $79.6$ $74.1$ $99.4$ $96.2$ $97.0$ $96.6$
Rl(Adapter) $0.44$ M 94.0 $69.4$ $79.7$ $74.2$ 99.4 $96.1$ $97.0$ $96.6$
Rl(LoRA) $0.39$ M $93.9$ $69.2$ $79.3$ $73.9$ 99.4 $96.2$ $97.0$ $96.6$
Rl(DoRA) $0.39$ M 94.0 $69.4$ 79.7 $74.2$ 99.4 $96.2$ 97.1 $96.6$
Rl(BiDoRA) $0.39$ M 94.0 71.3 $79.3$ 75.1 99.4 96.4 97.1 96.7
Further evidence of the effectiveness of BiDoRA can be observed in Table 4, which reports the results of token classification tasks. Unlike the NLU tasks discussed in the previous section, which involve classifying entire sentences and focusing on capturing global semantics, token classification requires classifying each token within a sentence, highlighting the importance of capturing local context. On the BioNLP dataset, BiDoRA consistently outperforms baseline methods by a large margin in terms of F1 score. On the CoNLL2003 dataset, BiDoRA either outperforms or matches all baseline methods across all metrics. Consistent with our previous findings, BiDoRA effectively fine-tunes pre-trained models for token classification tasks.

### v2

@@ 4.1 Overview of BiDoRA
Our method, BiDoRA, optimizes the trainable parameters in DoRA layers by solving a BLO problem. Let =$\mathcal{M}=\{\mathbf{m}_{1},\mathbf{m}_{2},\ldots,\mathbf{m}_{n}\}$ denote the set of magnitude components for all $n$ DoRA modules, and $\mathcal{V}=\{\Delta\mathbf{V}_{1},\Delta\mathbf{V}_{2},\ldots,\Delta\mathbf{V}_{n}\}$ denote the set of corresponding direction components. Specifically, we first learn the direction components $\mathcal{V}^{*}(\mathcal{M})$ on the training split of the downstream dataset $\mathcal{D}_{tr}$ at the lower level. The magnitude component $\mathcal{M}$ is tentatively fixed at this level; thus, the resulting optimal direction component $\mathcal{V}^{*}(\mathcal{M})$ is a function of $\mathcal{M}$. At the upper level, we determine the optimal magnitude component $\mathcal{M}^{*}$ by optimizing the loss on a validation split $\mathcal{D}_{val}$. In practice, $\mathcal{D}_{tr}$ and $\mathcal{D}_{val}$ are typically created by splitting the original training set without using additional data. This BLO problem is solved using an efficient gradient-based algorithm, where parameters at two levels are optimized iteratively until convergence. While this work focuses on the empirical validation of BiDoRA, our choice of optimization strategy is grounded in established theoretical research. The convergence properties of similar gradient-based bi-level algorithms have been previously analyzed (Pedregosa, 2016; Rajeswaran et al., 2019), providing confidence in the stability of our training procedure. Furthermore, the ability of such frameworks to improve generalization—a core objective of BiDoRA—has also been formally studied (Bao et al., 2021), supporting the rationale that our approach can mitigate overfitting.
@@ 4.2 Orthogonal Regularization
A central goal of BiDoRA is to learn the two disentangled components of a weight update: magnitude and direction. The direction component, $\Delta\mathbf{V}$, is responsible for finding a low-rank basis for the update directions. To maximize the expressive power of this component and prevent overfitting, its basis vectors (i.e., the columns of the direction matrix) should be as diverse and non-redundant as possible.
The orthogonality of neural network weights has been identified as a beneficial property (Bansal et al., 2018) and can effectively mitigate the overfitting issue (Balestriero & richard baraniuk, 2018). By enforcing orthogonality, the direction vectors are constrained to represent distinct, independent pathways for updates. This ensures that the limited parameter budget of the low-rank matrix is used efficiently to explore the solution space. Therefore, we define a Gram regularization loss (Xie et al., 2017) for the direction component:
$\mathcal{R}(\mathcal{V})=\sum_{k=1}^{n}\left\lVert(\mathbf{V}_{k}+\Delta\mathbf{V}_{k})^{\top}(\mathbf{V}_{k}+\Delta\mathbf{V}_{k})-\mathbf{I}\right\rVert_{F}^{2}$ (3)
where $\mathbf{I}$ is the identity matrix and $\|\cdot\|_{F}$ denotes the Frobenius norm. Intuitively, $\mathcal{R}(\mathcal{V})$ encourages each column of the direction matrix, representing a specific direction, to be orthogonal to one another. Since each column has already been normalized (equivalent to projected to the unit sphere), this also prompts each column to be far away from the other, thereby reducing the redundancy of parameters. The effectiveness of this constraint is empirically validated in our ablation study (See Table˜5), which shows a consistent performance improvement resulting from the enhanced generalization ability of the learned direction component.
@@ 4.3 A Bi-level Optimization Framework
Lower level.
At the lower level, we train the low-rank direction component $\mathcal{V}$ by minimizing a loss $\mathcal{L}_{tr}$ defined on the training set $\mathcal{D}_{tr}$. The overall training objective at this level is $\mathcal{L}_{tr}(\mathcal{V},\mathcal{M})=\mathcal{L}(\mathcal{V},\mathcal{M};\mathcal{D}_{tr})+\gamma\mathcal{R}(\mathcal{V})$. Here, $\mathcal{L}$ represents the fine-tuning loss, given the low-rank direction component $\mathcal{V}$, the magnitude component $\mathcal{M}$, and the training split $\mathcal{D}_{tr}$ of the downstream dataset. $\mathcal{R}(\mathcal{V})$ is the orthogonal regularizer defined in Eq.˜3, with $\gamma$ as a trade-off hyperparameter. In this level, we only update $\mathcal{V}$ while keeping $\mathcal{M}$ fixed, resulting in the following optimization problem:
@@
Upper level.
At the upper level, we validate the previously fixed magnitudes $\mathcal{M}$ on the validation set $\mathcal{D}_{val}$, using the optimal direction component $\mathcal{V}^{*}(\mathcal{M})$ that was learned at the lower level. This results in a validation loss $\mathcal{L}_{val}(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M})=\mathcal{L}(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M};\mathcal{D}_{val})$. We determine the optimal magnitude component $\mathcal{M}$ by minimizing this validation loss:
@@
A bi-level optimization framework.
Integrating the two levels of optimization problems, we have the following BLO framework:
@@
Two reasons exist behind the choice of setting the magnitude component as the upper level instead of the converse one: 1) In literature, the upper level usually has fewer parameters than the lower level. In our case, the design of setting the magnitude of complexity $\mathcal{O}(k)$ as the upper level and the direction of complexity $\mathcal{O}(dr+kr)$ as the lower level is consistent with the common practice. 2) BiDoRA resembles the DARTS method (Liu et al., 2019) in NAS literature, where the subnets are selected by a selection variable. Specifically, the magnitude vector resembles a selection variable on the direction matrix by softly selecting each direction (subnets) via scaling.
Optimization algorithm.
@@
2Initialize trainable magnitude components =$\mathcal{M}=\{\mathbf{m}_{k}\}_{k=1}^{n}$ and low-rank direction components $\mathcal{V}=\{\Delta\mathbf{V}_{k}\}_{k=1}^{n}=\{\{\mathbf{A}_{k}\}_{k=1}^{n},\{\mathbf{B}_{k}\}_{k=1}^{n}\}$
@@
We use a gradient-based optimization algorithm (Choe et al., 2023b) to solve the BLO problem presented in Section˜4.3. A significant challenge in this process is the computation of the upper-level loss gradient with respect to the magnitude component $\mathcal{M}$, as this gradient depends on the optimal solution of the lower-level problem, $\mathcal{V}^{*}(\mathcal{M})$. For deep neural networks, the lower-level objective is non-convex, meaning that finding the true optimal solution $\mathcal{V}^{*}(\mathcal{M})$ would require running its optimization process to full convergence. Performing this complete inner optimization for every single update of the upper-level variable $\mathcal{M}$ is computationally intractable.
To address this issue, we use the following one-step-unrolled approximation of $\mathcal{V}^{*}(\mathcal{M})$ inspired by previous work (Liu et al., 2019):
$\nabla_{\mathcal{M}}{\mathcal{L}_{val}(\mathcal{V}^{*}(\mathcal{M}),\mathcal{M})}\approx\nabla_{\mathcal{M}}{\mathcal{L}_{val}(\mathcal{V}-\xi\nabla_{\mathcal{V}}{\mathcal{L}_{tr}(\mathcal{V},\mathcal{M})},\mathcal{M})}$ (7)
@@
where $\epsilon$ is a small scalar and $\mathcal{V}^{\pm}=\mathcal{V}\pm\epsilon\nabla_{\bar{\mathcal{V}}}{\mathcal{L}_{val}(\bar{\mathcal{V}},\mathcal{M})}$. Since directly computing the matrix-vector multiplication term in Eq.˜8 is computationally expensive, we use finite difference to approximate this product as in Eq.˜9, following Liu et al. (2019). As detailed in Algorithm˜1, the direction component $\mathcal{V}$ and the magnitude component $\mathcal{M}$ are updated using gradient descent iteratively until convergence. After acquiring the optimal magnitudes $\mathcal{M}^{*}$ through the process above, the direction component $\mathcal{V}$ is retrained on the union of training and validation splits to achieve the best performance on downstream tasks, resulting in the final learned $\mathcal{V}^{*}$. All splits are intentionally used during retraining to maximize data utilization and performance.
In practice, the convergence of the search phase is determined by the evaluation metric at the upper level. For the subsequent retraining phase, we adopt a stopping criterion similar to DoRA’s, observing performance on a separate, held-out test set that is not used during training.
Table 1: RoBERTabase/large (Rb/l) and DeBERTaXXL (DXXL) with different fine-tuning methods on the GLUE benchmark (Wang et al., 2019). A higher value is better for all datasets. The best results are shown in bold.
Method #Params MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B Avg.
Rb(FT) 125.0125.0125.0M 90.390.390.3 94.894.894.8 89.389.389.3 61.661.661.6 86.786.786.7 92.892.892.8 76.976.976.9 91.291.291.2 85.585.585.5
Rb(Adapter) 0.90.90.9 M 86.586.586.5 94.094.094.0 88.488.488.4 58.858.858.8 92.592.592.5 89.189.189.1 71.271.271.2 89.989.989.9 83.883.883.8
Rb(LoRA) 0.150.150.15 M 86.886.886.8 94.394.394.3 88.088.088.0 60.360.360.3 93.0 89.689.689.6 72.972.972.9 90.190.190.1 84.484.484.4
Rb(DoRA) 0.170.170.17 M 86.886.886.8 94.294.294.2 89.289.289.2 60.560.560.5 92.992.992.9 89.689.689.6 73.273.273.2 90.2 84.684.684.6
Rb(BiDoRA) 0.170.170.17 M 87.1 94.4 89.4 61.3 92.792.792.7 90.6 76.0 90.190.190.1 85.2
Rl(FT) 355.0355.0355.0M 90.290.290.2 96.496.496.4 90.990.990.9 68.068.068.0 94.794.794.7 92.292.292.2 86.686.686.6 92.492.492.4 88.988.988.9
Rl(Adapter) 0.80.80.8M 90.390.390.3 96.396.396.3 87.787.787.7 66.366.366.3 94.7 91.591.591.5 72.972.972.9 91.591.591.5 86.486.486.4
Rl(LoRA) 0.390.390.39 M 90.6 96.396.396.3 90.090.090.0 66.966.966.9 94.594.594.5 91.291.291.2 86.386.386.3 91.791.791.7 88.488.488.4
Rl(DoRA) 0.390.390.39 M 90.6 96.4 89.889.889.8 65.865.865.8 94.7 91.291.291.2 86.686.686.6 92.0 88.488.488.4
Rl(BiDoRA) 0.390.390.39 M 90.6 96.196.196.1 90.1 67.0 94.694.694.6 91.7 86.9 92.0 88.6
DXXL(DoRA) 4.94.94.9M 91.2 96.3 92.3 71.1 95.3 91.6 91.8 90.8 90.0
DXXL(BiDoRA) 4.94.94.9M 91.7 96.3 92.6 72.3 95.2 92.0 92.3 90.8 90.4
@@ 5.1 Experimental Setup
We compare BiDoRA with several PEFT methods, including Full Fine-Tuning (FT), Adapter Tuning (Houlsby et al., 2019; Lin et al., 2020; Rücklé et al., 2021; Pfeiffer et al., 2021), LoRA (Hu et al., 2022a), A AdaLoRA (Zhang et al., 2023), DoRA (Liu et al., 2024a), VeRA (Kopiczko et al., 2024), FourierFT (Gao et al., 2024), AFLoRA (Liu et al., 2024b), LaMDA (Azizi et al., 2024), SSH (Shen et al., 2025b), MaCP (Shen et al., 2025a). BiDoRA does not use any additional data compared to other baselines, as we create the validation set for upper-level optimization by splitting the original training set with an 888:222 ratio for all tasks. Detailed descriptions of these baseline methods are provided in Appendix˜C.
Our experiments cover a wide range of tasks, including natural language understanding (Section˜5.2), extremely small biomedical datasets (Section˜5.3), natural language generation (Section˜H.1), and token classification (Section˜H.2). Please refer to the detailed dataset settings and experimental settings in Appendix˜A and Appendix˜B, respectively. Our implementation is based on the Huggingface Transformers library (Wolf et al., 2019) and the Betty library (Choe et al., 2023b).
@@ 5.2 Experiments on Natural Language Understanding Tasks
In this section, we evaluate the performance of BiDoRA on NLU tasks.
Main results.
@@
Table˜1 presents the results of fine-tuning the RoBERTa-base, RoBERTa-large, and DeBERTa XXL models on the GLUE benchmark with baseline PEFT methods and BiDoRA. The results show that BiDoRA achieves superior or comparable performance compared to baseline methods across all datasets with the same number of trainable parameters. The superior performance of BiDoRA verifies the effectiveness of its BLO mechanism. By training the magnitude and direction components on two distinct sub-datasets, BiDoRA enhances the flexibility of the learning process and improves learning capacity compared to DoRA, resulting in a performance boost.
We also present an experiment on the GLUE benchmark with the RoBERTa-base model, on a larger, wide range of baselines, following the settings from Shen et al. (2025b) and Shen et al. (2025a) and citing their reported baseline results for reference. The results in Table˜2 indicate that BiDoRA consistently outperforms all baselines, including DoRA, across these diverse NLU tasks, demonstrating its robust generalization capability.
Table 2: Performance of various fine-tuning methods on the GLUE benchmark for the RoBERTa-base model. The best ones are highlighted by bold and the second ones are highlighted by italic.
Model SST-2 MRPC CoLA QNLI RTE STS-B Avg.
FT 94.8 90.2 63.6 92.8 78.7 91.2 85.22
BitFit 93.7 92.7 62.0 91.8 81.5 90.8 85.42
AdptD 94.7 88.4 62.6 93.0 75.9 90.3 84.15
LoRA 95.1 89.7 63.4 93.3 78.4 91.5 85.23
AdaLoRA 94.5 88.7 62.0 93.1 81.0 90.5 84.97
AFLoRA 94.1 89.3 63.5 91.3 77.2 90.6 84.33
LaMDA 94.6 89.7 64.9 91.7 78.2 90.4 84.92
VeRA 94.6 89.5 65.6 91.8 78.7 90.7 85.15
FourierFT 94.2 90.0 63.8 92.2 79.1 90.8 85.02
SSH 94.1 91.2 63.6 92.4 80.5 90.9 85.46
MaCP 94.2 89.7 64.6 92.4 80.7 90.9 85.42
DoRA ($r=8$) 94.9 89.9 63.7 93.3 78.9 91.5 85.37
BiDoRA ($r=8$) 95.7 90.2 65.8 93.4 79.4 90.5 85.83
DoRA ($r=16$) 94.8 90.4 65.6 93.1 81.9 90.7 86.08
BiDoRA ($r=16$) 95.0 90.8 66.7 93.3 82.6 90.9 86.55
Table 3: Quantitative performance gap between training and test sets for DoRA and BiDoRA using the RoBERTa-base model. The gap is calculated as the training metric minus the test metric, where a smaller value indicates less overfitting.
Method SST-2 MRPC CoLA QNLI RTE STS-B Avg.
DoRA 2.0 9.5 32.5 6.6 18.0 8.8 12.9
BiDoRA 1.7 7.0 23.3 0.2 14.0 4.7 8.5
Robustness of BiDoRA towards different rank settings.
We explore the impact of different rank configurations on BiDoRA and DoRA, evaluating them with ranks of 888 and 161616 in addition to the rank of 444 used in Table˜1. The average accuracies reported in Table˜2 demonstrate that BiDoRA consistently surpasses DoRA across all rank configurations, highlighting its resilience and superior performance regardless of the rank setting.
Performance gap between training and testing set.
Figure 2: Training and test accuracy versus global training steps on the ModHayes split of the Reuters21578 dataset (Padmanabhan et al., 2016) when fine-tuning a RoBERTa-base model using DoRA and BiDoRA. The training and test curves for DoRA show a larger gap compared to BiDoRA, highlighting the effectiveness of our method in reducing overfitting.
[…]
5.3 Experiments on Extremely Small Datasets
Table 4: Fine-tuning ESM on the thermostability prediction task (Chen et al., 2023) (left), the BBP task (Dai et al., 2021) (middle), and the MIC task (Ledesma-Fernandez et al., 2023) (right). A higher value is better for all metrics except for MSE. The best results are highlighted in bold.
Methods #Params Accuracy Precision Recall F1
FT 652.7M 79.8 81.2 79.8 78.4
LoRA 1.5M 75.9 78.2 75.9 75.5
DoRA 1.6M 76.9 78.7 76.9 76.2
BiDoRA 1.6M 78.8 79.1 78.8 78.2
#Params Accuracy Precision Recall F1
652.9M 89.4 89.9 89.4 89.4
1.9M 86.8 87.7 86.8 86.7
2.0M 89.4 91.3 89.4 89.3
2.0M 92.1 93.1 92.1 92.0
#Params MSE
652.7M 0.2894
1.7M 0.3433
1.8M 0.2918
1.8M 0.2818
We conduct additional experiments on biomedical datasets, including two classification tasks—thermostability prediction (Chen et al. (2023), 936 training samples) and blood-brain barrier peptide prediction (BBP, Dai et al. (2021), 200 training samples)—and one regression task, minimum inhibitory concentration prediction (MIC, Ledesma-Fernandez et al. (2023), 3,695 training samples), which contain significantly fewer samples than standard NLP tasks.
The results are presented in Table˜4. Consistent with our previous findings, BiDoRA effectively fine-tunes pre-trained models on extremely small datasets. Our method outperforms the baselines by a larger margin as the dataset size decreases, confirming our previous conclusion that our method effectively combats the overfitting issue on various network architectures and diverse tasks.